# Bayesian Inference for Relational Graph in a Causal Vector Autoregressive Time Series

### v1

@@ 5.1 Simulation
For the simulation experiment, we primarily study the impact of the sample size and the sparsity level on estimation accuracy for the stationary precision matrix of a first order VAR process. We compare the accuracy of estimating the stationary precision matrix and the associated partial correlation graph for the proposed method along with two other popular graphical model estimation methods, the Gaussian Graphical Model (GGM) and the Gaussian Copula Graphical Model (GCGM).
@@
We generate the data from a VAR(1) model with a fixed marginal precision matrix, $\bm{\Omega}_{1}$. Specifically, the data $\bm{X}_{1},\ldots,\bm{X}_{T}$ are generated from the model
@@
with the stationary precision matrix chosen as $\bm{\Omega}=\bm{\Omega}_{1}$ and the parameters associated with the update of the conditional precision, $(\bm{L}_{1},\bm{K}_{1})$, are chosen as random $30\times 30$ with entries generated independently from $N(0,2.5^{2}).$ The VAR coefficient $\bm{A}_{1}$ and the innovation variance $\bm{\Sigma}_{1}$ are solved from the specified parameters, $\bm{\Omega}_{1},\bm{L}_{1},\bm{K}_{1},$ following the steps defined in section 3.1. The initial observation, $\bm{X}_{1}$, is generated from the stationary distribution and subsequent observations are generated via the iteration in (5.1). Two different sample sizes are used, $T=40$ and $T=60.$ For the sparse precision matrix, we used two different levels of sparsity, $15\%$ and $25\%$. The sparse precision $\bm{\Omega}_{1}$ is generated using the following method.
@@
    Adjacency matrix: we generate three small-world networks, each containing 10 disjoint sets of nodes with two different choices for nei variable in sample_smallworld of igraph (Csárdi et al., 2024) as 10 and 5. Then, we randomly connect some nodes across the small worlds with probability $q$. The parameters in the small world distribution and $q$ are adjusted to attain the desired sparsity levels in the precision matrix.
@@
    Precision matrix: Using the above adjacency matrix, we apply G-Wishart with scale 6 and truncate entries smaller than 1 in magnitude.
The simulated model is a full-rank 30-dimensional first-order VAR. Thus, the proposed method of fitting low-rank matrices for the conditional precision updates is merely an approximation and is fitting a possibly misspecified model. Posterior computation for the proposed method is done using the steps given in section 3.5, where a higher-order VAR with low-rank updates is used to fit the data. The upper bound on the order of the progress is chosen to be $p_{m}=10.$ Thus, we are fitting a mis-specified model, and evaluating the robustness characteristics of the method along with estimation accuracy. Table 1 shows the median MSE for estimation of the 30$\times$30 precision matrix for model (5.1). The proposed method has higher estimation accuracy than the GCGM and GGM methods that ignore the temporal dependence. Thus, explicitly modeling the dependence, even under incorrect specification of the rank, provides substantial gain in terms of MSE. The comparative performance is better for the proposed method when the sparsity level is $25\%.$ The Gaussian Graphical Model seemed sensitive to the specification of the simulation model. The MSE for he GGM increases with increasing sample size, a phenomenon presumably an artifact of ignoring the dependence in the sample. We simulated the parameters of the conditional precision update, $\bm{L}_{1},\bm{K}_{1},$ independently from $N(0,1)$ (not reported here) as well and found that the MSE for the GGM to be decreasing with the sample size. This could be because the $N(0,1)$ is more concentrated around zero, making the simulated model closer to the independent model. The MSE for all the methods goes up with a decreased sparsity level. A greater number of nonzero entries in the precision matrix makes the problem harder for sparse estimation methods and leads to higher estimation errors. Our results above are all based on rank-1 updates. We also obtain results for rank-3 updates (not shown here) but the results show marginal improvement over rank-1 updates for precision matrix estimation.
We also investigate how well the partial correlation graph is estimated under different methods when the samples are dependent, arising from a VAR process. The true graph is given by the nonzero entries in the true precision $\bm{\Omega}$. The estimated graph is obtained by thresholding the estimated precision at a given level $\tau.$ For the present investigation we use $\tau=0.15$, i.e. two different time series nodes are declared connected if the partial correlation between the two components exceeds 0.15 in the estimated precision matrix. Figure 1 shows the ROC curves for the three estimation methods for different settings; two different sample sizes $T=40,60$ and two different sparsity levels $15\%,25\%.$ In every case the proposed method has better performance than GCGM and GGM, with the GCGM performing better than the GGM. The AUC values for the proposed method are all bigger than those for the GCGM and GGM.
@@
5.2 Graph Structure in US GDP Components
For graph estimation, the simulation experiment illustrated the benefit of explicitly accounting for temporal dependence in the sample. Here we analyze the graphical association among components of the US GDP based on time series data on each of the components. The data are obtained from bea.gov, collected quarterly for the period 2010 to 2019 with a total of 40 time points. We chose this particular period to avoid large external shocks like the Great Recession and the COVID-19 pandemic. We study 14 components that are used in the computation of the aggregate GDP value. Specifically, we study associations between the following components: Durable Goods (Dura), Nondurable Goods (NonDur), Services (Serv), Structures (Strct), Equipment (Equip), Intellectual Property Products (IPP), Residential Products (Resid), Exports-Goods (Exp-G), Exports-Services (Exp-S), Imports-Goods (Imp-G), Imports-Services (Imp-S), National Defense (NatDef), Nondefense expenditure (NonDef), State and Local expenditures (St-Lo). The time series for the components are reported independently. However, several of these variables share implicit dependencies. We apply our proposed model to study these dependencies.
The graphical structure is established based on the partial correlation graph obtained from estimating the stationary precision matrix. Figure 2 illustrates the estimated graphical dependence based on the estimated partial correlations with an absolute value of more than 0.15. Thresholding the partial correlations at 0.15 to determine active edges leads to 18 connections, about a fifth of the total possible edges. The partial correlation threshold was varied to check the sensitivity of the estimated graph to the choice of the threshold over the values ranging from 0.05 to 0.5.
Figure 2: The estimated graphical association among 14 macro-economic variables governing US GDP (threshold 0.15).
Table 2: Connectivity by thresholding partial correlations at 0.15; connections are denoted by ‘*’.
Dura Nondur Serv Strct Equip IPP Resid Exp-G Exp-S Imp-G Imp-S NatDef NonDef St-Lo
Dura * *
NonDur *
Serv *
Strct * *
Equip * * * *
IPP * *
Resid * * * * *
Exp-G * * * * *
Exp-S * * * * *
Imp-G *
Imp-S * * *
NatDef * *
NonDef * *
St-Lo *
To infer the estimated network, we compute graph theoretic nodal attributes. Specifically, we consider 1) Betweenness centrality, 2) Node impact, 3) Degree, 4) Participation coefficient, 5) Efficiency, 6) Average shortest path length, 7) Shortest path eccentricity, and 8) Leverage. For completeness, we provide a short description of these attributes in Section X of the supplementary materials.
From the attributes for the 14 GDP components, equipment, residential products, and export/import goods and services show a high degree of betweenness and centrality. These nodes seem to be deep in the graph. A better picture would emerge if the GDP components are studied at a more granular level in terms of their subcomponents. We plan to investigate that in the future.
Table 3: Graph attributes, computed based on the graphical association in Figure 2.
Dura Nondur Serv Strct Equip IPP Resid Exp-G Exp-S Imp-G Imp-S NatDef NonDef St-Lo
Betweenness 9.000 0.000 0.000 10.000 39.000 10.000 34.000 50.000 34.000 0.000 14.000 2.000 5.000 0.000
Node impact 0.027 -0.101 -0.062 0.066 -0.104 0.040 -0.201 -0.577 0.425 -0.101 0.168 -0.075 -0.024 -0.101
Degree 2.000 1.000 1.000 2.000 4.000 2.000 5.000 5.000 5.000 1.000 3.000 2.000 2.000 1.000
Participation coefficient 0.500 1.000 1.000 1.000 0.375 0.500 0.440 0.440 0.440 1.000 0.556 1.000 1.000 1.000
Efficiency 0.000 0.000 0.000 0.000 0.000 0.000 0.100 0.000 0.100 0.000 0.333 0.000 0.000 0.000
Avg shortest pathlen 2.071 2.786 2.571 2.071 1.714 2.000 1.929 1.929 1.643 2.786 1.857 2.643 2.357 2.786
Shortest path eccentricity 3.000 5.000 4.000 3.000 3.000 3.000 4.000 4.000 3.000 5.000 3.000 5.000 4.000 5.000
Leverage -0.381 -0.667 -0.600 -0.314 0.178 -0.429 0.355 0.460 0.244 -0.667 -0.100 -0.214 -0.214 -0.667
@@ 6 Discussion
In this article, we propose a new Bayesian method for estimating the stationary precision matrix of a high-dimensional VAR process under stability constraints and subsequently estimating the contemporaneous stationary graph for the components of the VAR process. The method has several natural advantages. The methodology introduces a parameterization that allows fast computation of the stationary likelihood of a reduced rank high-dimensional Gaussian VAR process, a popular high-dimensional time series model. The new parameterization introduced here gives a natural way to directly model the sparse stationary precision matrix of the high-dimensional VAR process, which is the quantity needed to construct contemporaneous stationary graphs for the VAR time series. Most estimation methods for high-dimensional VAR processes fail to impose causality (or even stationarity) on the solution, thereby making estimating stationary precision matrix less meaningful. The proposed methodology uses causality as a hard constraint so that any estimate of the VAR process is restricted to the causal VAR space. We also show the posterior consistency of our Bayesian estimation scheme when priors are defined through the proposed parameterization. The focus of the present article is on the estimation of the stationary precision matrix of a high-dimensional stationary VAR. Hence, the direct parameterization of the precision, independent from other parameters involved in the temporal dynamics, is critical. However, a similar scheme can provide a direct parameterization of the VAR autocovariance matrices and, hence, the VAR coefficients under reduced rank and causality constraints for a high-dimensional VAR. This is part of future investigation. We also plan to investigate graph consistency for the partial correlation graph obtained by thresholding the entries of the stationary precision matrix of a high-dimensional VAR.
@@
We follow arguments in the general posterior contraction rate result from Theorem 8.19 of Ghosal and van der Vaart [2017] on the joint density of the entire multivariate time series by directly constructing a likelihood ratio test satisfying a condition like (8.17) of Ghosal and van der Vaart [2017]. We also use the simplified prior concentration condition (8.22) and global entropy condition (8.23) of Ghosal and van der Vaart [2017]. It is more convenient to give direct arguments than to fit within the notations of the theorem.
@@ A.1 Prior concentration: pre-rate
We note that by Relation (iii) of Lemma 5, $\|\bm{\Upsilon}_{T,0}\|^{2}_{\mathrm{op}}\leq\|\bm{\Sigma}_{0}\|^{2}_{\mathrm{op}}$, while by Relation (i), $\|\bm{\Upsilon}_{T,0}^{-1}\|^{2}_{\mathrm{op}}\leq\|\bm{\Sigma}_{0}^{-1}\|^{2}_{\mathrm{op}}$, which are bounded by a constant by Condition (A5). Next observe that by Relation (iii) applied to $\bm{\Upsilon}_{T}$, $\|\bm{\Upsilon}_{T}\|_{\mathrm{op}}\leq\|\bm{\Sigma}\|_{\mathrm{op}}\leq\|\bm{\Omega}^{-1}\|_{\mathrm{op}}=\|1/\bm{f}\|_{\infty}$, where $1/\bm{f}$ refers to the vector of the reciprocals of the entries of $\bm{f}$. Since the entries of $\bm{f}_{0}$ lie between two fixed positive numbers, so do the entries of $\bm{f}$ and $1/\bm{f}$ when $\|\bm{f}-\bm{f}_{0}\|_{\infty}$ is small. Then it is immediate that (see, e.g., Lemma 4 of Roy et al. [2024]) that $\|\bm{\Omega}-\bm{\Omega}_{0}\|_{\mathrm{F}}\lesssim d\eta$. Applying Lemma 8, we obtain the relation $\|\bm{\Upsilon}_{T,0}-\bm{\Upsilon}_{T}\|_{\mathrm{F}}\lesssim drT^{2}\eta$, which is small for $\eta$ sufficiently small. Thus, $\|\bm{\Upsilon}_{T,0}^{-1/2}(\bm{\Upsilon}_{T,0}-\bm{\Upsilon}_{T})\bm{\Upsilon}_{T,0}^{-1/2}\|_{\mathrm{F}}$ is also small. Since the operator norm is weaker than the Frobenius norm, this implies all eigenvalues of $\bm{\Upsilon}_{T,0}^{-1/2}(\bm{\Upsilon}_{T,0}-\bm{\Upsilon}_{T})\bm{\Upsilon}_{T,0}^{-1/2}$ are close to zero. Now, expanding the estimates in Parts (ii) and (iii) of Lemma 4 in a Taylor series, the Kullback-Leibler divergence and variation between $Q_{\bm{\zeta}_{0}}$ and $Q_{\bm{\zeta}}$ are bounded by a multiple of $\|\bm{\Upsilon}_{T,0}^{-1/2}(\bm{\Upsilon}_{T,0}-\bm{\Upsilon}_{T})\bm{\Upsilon}_{T,0}^{-1/2}\|_{\mathrm{F}}^{2}\leq d^{2}r^{2}T^{4}\eta^{2}=T\bar{\epsilon}_{T}^{2}$ if $\eta$ is chosen to be $d^{-1}T^{-3/2}\epsilon_{T}$. The prior probability of $\|\bm{\zeta}-\bm{\zeta}_{0}\|_{\infty}\leq\eta$ is, in view of the assumed a priori independence of all components of $\bm{\zeta}$, bounded below by $(\bar{c}\eta)^{\mathrm{dim}(\bm{\zeta})}$ for some constant $\bar{c}>0$. Hence $-\log\Pi(\|\bm{\zeta}-\bm{\zeta}_{0}\|_{\infty}\leq\eta)\lesssim(d^{2}+d+2pdr)\log(1/\eta)\lesssim d^{2}\log(1/\eta)$. Thus (A.1) for $T\bar{\epsilon}_{T}^{2}\asymp d^{2}\log T$, that is for $\bar{\epsilon}_{T}\asymp d\sqrt{(\log T)/T}$, or any larger sequence.

### v2

@@ 7 Simulation
To analyze the performance of the proposed method when the ground truth is known, we perform further simulation experiments. We primarily study the impact of the sample size and the sparsity level on estimation accuracy for the stationary precision matrix for a first order VAR process. We compare the accuracy of estimating the stationary precision matrix and the associated partial correlation graph for the proposed method with two other popular graphical model estimation methods, the Gaussian Graphical Model (GGM) and the Gaussian Copula Graphical Model (GCGM).
@@
The data are generated from a VAR(1) model with a fixed marginal precision matrix, $\bm{\Omega}_{1}$. Specifically, the data $\bm{X}_{1},\ldots,\bm{X}_{T}$ are generated from the model
@@
with the stationary precision matrix chosen as $\bm{\Omega}=\bm{\Omega}_{1}$ and the parameters associated with the update of the conditional precision, $(\bm{L}_{1},\bm{K}_{1})$, are chosen as random $30\times 30$ matrices with entries generated independently from $\mathrm{N}(0,2.5^{2}).$ The VAR coefficient $\bm{A}_{1}$ and the innovation variance $\bm{\Sigma}_{1}$ are solved from the specified parameters, $\bm{\Omega}_{1},\bm{L}_{1},\bm{K}_{1},$ following the steps defined in section 4.1. The initial observation, $\bm{X}_{1}$, is generated from the stationary distribution, and subsequent observations are generated via the iteration in (7.1). Two different sample sizes are used, $T=40$ and $T=60.$ For the sparse precision matrix, we use two different levels of sparsity, $15\%$ and $25\%$. The sparse precision $\bm{\Omega}_{1}$ is generated using the following method.
@@
    Adjacency matrix: three small-world networks are generated, each containing 10 disjoint sets of nodes with two different choices, 5 and 10, for nei variable in sample_smallworld of igraph (Csárdi et al., 2024). Then, nodes are randomly connected across the small worlds with probability $q$. The parameters in the small world distribution and $q$ are adjusted to attain the desired sparsity levels in the precision matrix.
@@
    Precision matrix: Using the above adjacency matrix, the precision matrix $\bm{\Omega}$ is generated from a G-Wishart with scale equal to 6, and off-diagonal entries smaller than 1 in magnitude are set to zero.
The simulated model is a full-rank 30-dimensional first-order VAR. Thus, the proposed method of fitting low-rank matrices for the conditional precision updates is merely an approximation and is fitting a possibly misspecified model. Posterior computation for the proposed method uses the steps in section 4.4, where a higher-order VAR with low-rank updates is used to fit the data. The upper bound on the order of the progress is chosen to be $p_{m}=10.$ Thus, we fit a misspecified model and evaluate the robustness characteristics of the method along with estimation accuracy. Table 3 shows the median MSE for estimating the entries of the 30$\times$30 precision matrix for model (7.1). The proposed method has higher estimation accuracy than the GCGM and GGM methods that ignore the temporal dependence. Explicit modeling of the dependence, even under incorrect rank specification, is beneficial and can provide substantial gains in MSE. The comparative performance is better for the proposed method when the sparsity level is $25\%.$ The Gaussian Graphical Model seems sensitive to the specification of the simulation model. The MSE for the GGM increases with increasing sample size, a phenomenon presumably an artifact of ignoring the dependence in the sample. When the simulated parameters of the conditional precision update, $\bm{L}_{1},\bm{K}_{1},$ are generated independently from $\mathrm{N}(0,1)$ (not reported here), the MSE for the GGM decreases monotonically with the sample size. This is because $\mathrm{N}(0,1)$ being more concentrated around zero, makes the simulated model closely mimic an independent model. The MSEs for all the methods go up with a decreased sparsity level. A higher number of nonzero entries in the precision matrix makes the problem harder for sparse estimation methods and leads to higher estimation errors. The results for the proposed method presented in this section are based on rank-1 updates. The results for rank-3 updates (not shown here) show only marginal improvement over rank-1 updates for precision matrix estimation. The rank-3 updates are computationally more expensive.
We also investigate how well the partial correlation graph is estimated under different methods when the samples are dependent and arise from a VAR process. Figure 3 shows the ROC curves for the three estimation methods for different settings; two different sample sizes $T=40,60$ and two different sparsity levels $15\%,25\%.$ In every case, the proposed method performs better than GCGM and GGM, with the GCGM performing better than the GGM. The AUC values for the proposed method are bigger than those for the GCGM and GGM.
@@ 8 Discussion
In this article, we propose a new Bayesian method for estimating the stationary precision matrix of a high-dimensional VAR process under stability constraints and subsequently estimating the contemporaneous stationary graph for the components of the VAR process. The method is then used to evaluate possible changes during the financial crisis of the Great Recession of 2007–2009 in terms of the interdependencies among various seasonally adjusted quarterly time series collected in the FRED-QD database. The method has several natural advantages that will be useful in analyzing similar time series databases. The methodology introduces a parameterization that allows for the fast computation of the stationary likelihood of a reduced-rank high-dimensional Gaussian VAR process. The new parameterization also gives a natural way to directly model the sparse stationary precision matrix of a high-dimensional VAR process, which is the quantity needed to construct contemporaneous stationary graphs for the VAR time series. Another distinct advantage of the proposed method over popular competitors is that the proposed methodology uses causality of the VAR process as a hard constraint so that any estimated VAR will be restricted to the causal VAR space. This property is helpful in applications, such as stable dynamical system identification, where stationarity of the process is paramount. Most estimation methods for high-dimensional VAR processes fail to impose causality on the solution, making estimating the stationary precision matrix less meaningful. In our preliminary analysis of the FRED-QD data using sparse VAR, the estimated VAR polynomials are often unstable; the estimated VAR process is non-causal. However, estimates based on the proposed parameterization are always stable by construction.
We also establish posterior consistency for our Bayesian estimation scheme when priors are defined through the proposed parameterization. The consistency of $\bm{\Omega}$ posterior indicates accurate inference on change parameters $\theta_{g}(i,j)$ to determine the effect of the Great Recession on the individual edges. In the future, we plan to investigate graph consistency for the partial correlation graph obtained by thresholding the entries of the stationary precision matrix of a high-dimensional VAR.
The focus of the present article is on the estimation of the stationary precision matrix of a high-dimensional stationary VAR. Hence, the direct parameterization of the precision, independent from other parameters involved in the temporal dynamics, is critical. However, a similar scheme can provide a direct parameterization of the VAR autocovariance matrices and, hence, the VAR coefficients under reduced rank and causality constraints for a high-dimensional VAR. This is part of a future investigation. Another possibility is to estimate directional relationships among the variables in each group of the FRED-QD data through directed acyclic graph (DAG)-based characterization.
Finally, the methodology developed here was motivated by the desire to estimate and compare the stationary graphs between the economic indicators before and after the Great Recession. But the methods are broadly applicable for doing a similar before-and-after analysis of VAR time series graphs in the context of other extreme events, such as the COVID-19 pandemic, provided the event is a prolonged and complex event that would be hard to analyze using change point analysis. In addition, the methods apply to any problem where the main interest is the estimation of the partial correlation graph of a VAR process.
@@
Appendix: Identification of Parameters
In the recursive algorithm described above, the parameters $\bm{\Omega},(\bm{L}_{1},\bm{K}_{1}),\ldots,(\bm{L}_{p},\bm{K}_{p})$ are mapped to the modified parameters $\bm{\Omega},$ $(\bm{U}_{1},\bm{V}_{1}),\ldots,(\bm{U}_{p},\bm{V}_{p})$. The mapping is not one-to-one since each $\bm{V}_{j}$ has the restriction $\bm{V}_{j}^{\mathrm{T}}\bm{D}_{j-1}^{-1}\bm{V}_{j}=\bm{I}$ while the basic parameters $\bm{K}_{j}$ do not have any restrictions. The intermediate parameters in the mapping, $\bm{\Omega},(\bm{U}_{1},\bm{V}_{1}),\ldots,(\bm{U}_{p},\bm{V}_{p})$ are identifiable only up to rotations. Consider the equivalence classes of pair of $d\times r_{j}$ matrices $(\bm{U},\bm{V})$ defined through the relation $(\bm{U},\bm{V})\equiv(\bm{U}\bm{Q},\bm{V}\bm{Q})$ for any $r_{j}\times r_{j}$ orthogonal matrix $\bm{Q}$. Let for each $r_{j}$, $\mathbb{C}_{r_{j}}$ be the set of such equivalence classes. Also let $\mathcal{S}^{++}_{d,s}$ be a subset of the positive definite cone $\mathcal{S}^{++}_{d}$ with specified sparsity level $s$. The following proposition shows that there is a bijection from the set $\mathcal{S}^{++}_{d,s}\times\mathbb{C}_{r_{1}}\times\cdots\times\mathbb{C}_{r_{p}}$ to the causal VAR$(p)$ parameter space defined by the VAR coefficients $\bm{A}_{1},\ldots,\bm{A}_{p}$ and innovation variance $\bm{\Sigma}$, where polynomials with coefficients $\bm{A}_{1},\ldots,\bm{A}_{p}$ satisfy low-rank restrictions and the stationary precision satisfies $s$-sparsity. The definition of sparsity is made clearer in the prior specification section. For simplicity, we use a mean-zero VAR($p$) process. Before formally stating the identification result, we state and prove a short lemma used to show identification.
Lemma 3.
Given an $n\times n$ matrix $\bm{A}$ of rank $1\leq r\leq n$ for some positive integer $n$, and a positive definite $n\times n$ matrix $\bm{B}$, there exists $n\times r$ matrices $\bm{C}$ and $\bm{R}$ such that $\bm{A}=\bm{C}\bm{R}^{\mathrm{T}}$, $\mathrm{rank}(\bm{C})=\mathrm{rank}(\bm{R})=r$ and $\bm{R}^{\mathrm{T}}\bm{B}\bm{R}=\bm{I}$.
Proof.
Given a rank factorization $\bm{A}=\bm{U}\bm{V}$ and the symmetric square root $\bm{B}^{1/2}$, let $\bm{S}=\bm{Q}\bm{Z}$ be the Q-R decomposition of $\bm{S}=\bm{B}^{1/2}\bm{V}^{\mathrm{T}}$ where $\bm{Q}$ is $n\times r$ semi-orthogonal and $\bm{Z}$ is $r\times r$ nonsingular matrix. Then
$\bm{A}=\bm{U}\bm{V}=\bm{U}\bm{V}\bm{B}^{1/2}\bm{B}^{-1/2}=\bm{U}\bm{V}\bm{B}^{1/2}\bm{B}^{-1/2}=\bm{U}\bm{Z}^{\mathrm{T}}\bm{Q}^{\mathrm{T}}\bm{B}^{-1/2}=\bm{C}\bm{R},$
where $\bm{C}=\bm{U}\bm{Z}^{\mathrm{T}}$ and $\bm{R}=\bm{B}^{-1/2}\bm{Q}$. For this choice, $\bm{R}^{\mathrm{T}}\bm{B}\bm{R}=\bm{I}.$
Proposition 2.
Let $\bm{\Omega}>0$ be a $d\times d$ positive definite matrix, which is $s$-sparse. Let $(\bm{U}_{1},\bm{V}_{1}),\ldots$, $(\bm{U}_{p},\bm{V}_{p})$ be given full column rank matrix pairs, of order $d\times r_{1},\dots,d\times r_{p}$, respectively. Then there is a unique causal $\mathrm{VAR}(p)$ process with stationary variance matrix $\bm{\Gamma}(0)=\bm{\Omega}^{-1}$ and autcovariances $\bm{\Gamma}(1),\ldots,\bm{\Gamma}(p)$ (and hence the coefficients $\bm{A}_{1},\ldots,\bm{A}_{p}$ as in the Yule-Walker solution) uniquely determined recursively from $\bm{W}_{j}=\bm{U}_{j}\bm{V}_{j}^{\mathrm{T}}$. The associated increments in the conditional precision matrices $\bm{C}_{j}^{-1}-\bm{C}_{j-1}^{-1}$ will be of rank $r_{j}.$
Proof.
@@
We follow arguments in the general posterior contraction rate result from Theorem 8.19 of Ghosal and van der Vaart (2017) on the joint density of the entire multivariate time series by directly constructing a likelihood ratio test satisfying a condition like (8.17) of Ghosal and van der Vaart (2017). We also use the simplified prior concentration condition (8.22) and global entropy condition (8.23) of Ghosal and van der Vaart (2017). It is more convenient to give direct arguments than to fit within the notations of the theorem.
@@ A.1 Prior concentration: pre-rate
We note that by Relation (iii) of Lemma 5, $\|\bm{\Upsilon}_{T,0}\|^{2}_{\mathrm{op}}\leq\|\bm{\Sigma}_{0}\|^{2}_{\mathrm{op}}$, while by Relation (i), $\|\bm{\Upsilon}_{T,0}^{-1}\|^{2}_{\mathrm{op}}\leq\|\bm{\Sigma}_{0}^{-1}\|^{2}_{\mathrm{op}}$, which are bounded by a constant by Condition (A5). Next observe that by Relation (iii) applied to $\bm{\Upsilon}_{T}$, $\|\bm{\Upsilon}_{T}\|_{\mathrm{op}}\leq\|\bm{\Sigma}\|_{\mathrm{op}}\leq\|\bm{\Omega}^{-1}\|_{\mathrm{op}}=\|1/\bm{f}\|_{\infty}$, where $1/\bm{f}$ refers to the vector of the reciprocals of the entries of $\bm{f}$. Since the entries of $\bm{f}_{0}$ lie between two fixed positive numbers, so do the entries of $\bm{f}$ and $1/\bm{f}$ when $\|\bm{f}-\bm{f}_{0}\|_{\infty}$ is small. Then it is immediate that (see, e.g., Lemma 4 of Roy et al. (2024)) that $\|\bm{\Omega}-\bm{\Omega}_{0}\|_{\mathrm{F}}\lesssim d\eta$. Applying Lemma 8, we obtain the relation $\|\bm{\Upsilon}_{T,0}-\bm{\Upsilon}_{T}\|_{\mathrm{F}}\lesssim drT^{2}\eta$, which is small for $\eta$ sufficiently small. Thus, $\|\bm{\Upsilon}_{T,0}^{-1/2}(\bm{\Upsilon}_{T,0}-\bm{\Upsilon}_{T})\bm{\Upsilon}_{T,0}^{-1/2}\|_{\mathrm{F}}$ is also small. Since the operator norm is weaker than the Frobenius norm, this implies all eigenvalues of $\bm{\Upsilon}_{T,0}^{-1/2}(\bm{\Upsilon}_{T,0}-\bm{\Upsilon}_{T})\bm{\Upsilon}_{T,0}^{-1/2}$ are close to zero. Now, expanding the estimates in Parts (ii) and (iii) of Lemma 4 in a Taylor series, the Kullback-Leibler divergence and variation between $Q_{\bm{\zeta}_{0}}$ and $Q_{\bm{\zeta}}$ are bounded by a multiple of $\|\bm{\Upsilon}_{T,0}^{-1/2}(\bm{\Upsilon}_{T,0}-\bm{\Upsilon}_{T})\bm{\Upsilon}_{T,0}^{-1/2}\|_{\mathrm{F}}^{2}\leq d^{2}r^{2}T^{4}\eta^{2}=T\bar{\epsilon}_{T}^{2}$ if $\eta$ is chosen to be $d^{-1}T^{-3/2}\epsilon_{T}$. The prior probability of $\|\bm{\zeta}-\bm{\zeta}_{0}\|_{\infty}\leq\eta$ is, in view of the assumed a priori independence of all components of $\bm{\zeta}$, bounded below by $(\bar{c}\eta)^{\mathrm{dim}(\bm{\zeta})}$ for some constant $\bar{c}>0$. Hence $-\log\Pi(\|\bm{\zeta}-\bm{\zeta}_{0}\|_{\infty}\leq\eta)\lesssim(d^{2}+d+2pdr)\log(1/\eta)\lesssim d^{2}\log(1/\eta)$. Thus (A.1) for $T\bar{\epsilon}_{T}^{2}\asymp d^{2}\log T$, that is for $\bar{\epsilon}_{T}\asymp d\sqrt{(\log T)/T}$, or any larger sequence.